Graham, a physics enthusiast, is asked by his instructor to visit a nearby carpentry. He observed
there a hand–held powered reciprocating saw, say a sabre saw, which performs SHM in a vertical plane with
an angular frequency of 350 rad s−1
. Assume that the tip of the toothed blade of the saw is at x = 0 with
positive velocity when t = 0 and the amplitude of SHM is 2.40 cm. Write an equation, as a function of time,
which shows the position of the tip of the toothed blade. Also calculate that how much will the blade take to
travel from x = 0 to x = 1.20 cm? To 2.40 cm?
1
Expert's answer
2019-10-28T11:03:07-0400

Let the position of the tip of the toothed blade be represented by the SHM equation "x=A\ sin(\omega t+\phi)" ............(1)

where "x" is the displacement of the tip at time "t" , "A" is the amplitude, "\omega" is the angular frequency, and "\phi" is the initial phase.

Given "\omega=350\ rad/s" , "A=2.40\ cm"

Is it also given that at "t=0," "x=0"

Therefore, from (1),

"0=A\ sin(0+\phi)\\ \Rightarrow sin\phi=0\\ \Rightarrow \phi=0"

"\therefore" From (1), "x=A\ sin(\omega t+0)"

"\Rightarrow x=A\ sin(\omega t)"

Velocity, "v=\frac{dx}{dt}=A\omega\ cos(\omega t)"

At "t=0" , "v=A\omega" , which is positive and consistent with the problem.

So, "x=A\ sin(\omega t)\Rightarrow \boxed {x=2.40 cm\ sin(350 t)}"

At "x=1.20\ cm"

"1.20\ cm=2.40\ cm\ sin(350 t)"

"sin(350t)=0.5\Rightarrow sin(350t)=sin(\pi/6)"

"\Rightarrow 350t=\pi/6\Rightarrow t=\frac{\pi}{6\times 350}\ seconds"

"t=0.00150\ seconds = 1.50 \ ms"

At "x=2.40\ cm" ,

"2.40\ cm=2.40\ cm\ sin(350t)\\
sin(350t)=1\\
sin(350t)=sin(\pi/2)\\
350t=\pi/2\\
t=\frac{\pi}{350\times 2}\\
t=0.00449\ seconds = 4.49\ ms"


Answer: The position of the tip at any time is given by "x=2.40cm\ sin(350t)"

Time taken to travel from x=0 to x=1.20 cm is 1.50 ms.

Time taken to travel from x=0 to x=2.40 cm is 4.49 ms